OpenStudy (anonymous):

WILL AWARD MEDALS Find the rectangular coordinates of the point with the polar coordinates (1, 1/2 pi)

OpenStudy (anonymous):

\[(1, \frac{ 1 }{ 2 } \pi )\]

OpenStudy (anonymous):

\[x=r\cos(\theta), y=r\sin(\theta)\] cmpute \[x=\cos(\frac{\pi}{2}),y=\sin(\frac{\pi}{2})\]

OpenStudy (anonymous):

?? a. (0, -1) b. (1, 0) c. (0, 1) d. (-1, 0)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Well the angle is measured counter clockwise from the right, or EAST direction. So \[\pi/2\] radians is 90 degrees. If you go 90 degrees counter clockwise from the right which one (x or y) will be 1 and which will be 0 ?

OpenStudy (anonymous):

(0,1)?

OpenStudy (anonymous):

@Dom89

OpenStudy (anonymous):

yepyep